OpenStudy (anonymous):

1,440 is 120 percent of

OpenStudy (anonymous):

1200

OpenStudy (anonymous):

1440 = 120% * x 1440 = 1.2 * x x = 1440/1.2 x = 1200

OpenStudy (anonymous):

let the no be = x 120 1440 * 100 ---- x = 1440 => x = ----------- = 1200 100 120